Circular No. 5988

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994O IN MARKARIAN 268 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova of mag 18–19 in Markarian 268 on films taken May 6.93 and 7.88 UT. SN 1994O is located at α = 13^{h}38^{m}53^{s}.85, δ = +30^{o}37′51′′.5 (equinox 1950.0), which is superimposed on the dark central part of the galaxy and is 2′′.5 west and 2′′.5 north of the galaxy’s center. Nothing is visible at this location on films taken Apr. 8.99 and earlier. A nearby star of V ∼ 15 has end figures 57^{s}.01, 36′33′′.8.

P. Challis, Harvard-Smithsonian Center for Astrophysics; and L. Wells, National Optical Astronomy Observatory, report their confirmation of SN 1994O from imaging obtained at the 1.2-m telescope at Mt. Hopkins on May 11 and 12 UT. A spectrogram (range 320–800 nm) obtained at the Multiple Mirror Telescope (+ Blue Channel) on May 13 shows strong P-Cyg Si II (635.5 nm), indicating that this object is a type-Ia supernova within a week of maximum light. The redshift of the host galaxy, derived from the Ca H and K absorption, is cz ∼ 0.04.

PERIODIC COMET BROOKS 2 (1994j) S. Nakano, Sumoto, Japan, reports the recovery of this comet indepen-

dently by A. Nakamura and T. Seki, the following observations showing a correction of ΔT = −0.01 day to the prediction on MPC 18259 (ephemeris on MPC 23143):

1994 UT α_{2000} δ_{2000} m_{1} Observer
May 8.77813 23^{h}14^{m}45.^{s}40 −2^{o}03′00′′.8 18.2 Nakamura
8.78524 23 14 46.27 −2 02 56.3  _{”}
8.78924 23 14 46.54 −2 02 52.8  _{”}
12.77882 23 22 50.06 −1 16 08.7 19  Seki

A. Nakamura (Kuma Kogen). 0.60-m f/6.0 Ritchey-Chrétien + CCD. T. Seki (Geisei). 0.60-m f/3.5 reflector. Comet condensed with coma di-

ameter 7′′.

PERIODIC COMET SCHWASSMANN-WACHMANN 1 Total V magnitude estimates from H. Mikuž, Ljubljana, Slovenia (0.20-

m Baker-Schmidt camera + CCD): Mar. 30.86 UT, 13.6; Apr. 7.85, 13.7; 29.84, 12.9; May 3.83, 12.7; 9.85, 12.7.

1994 May 14 Daniel W. E. Green